OpenStudy (anonymous):

Find c such that f_ave=f(c) of 6sin(x)-3sin(2x).

OpenStudy (anonymous):

I have found the average value of the function to be 12/pi. Basically, I just need help solving the equation for c. I'm stuck and I've been trying for awhile now. Here are the steps I've taken so far: \[6\sin(c)-3\sin(2c)=12/\pi\] \[6\sin(c)-3(\sin(c)\cos(c)+\sin(c)\cos(c))\] \[6\sin(c)(1-\cos(c))\] \[\sin(c)(1-\cos(c))=2/\pi\] It's at this point that I get stuck. Am I missing something?